[Date Prev][Date Next][Thread Prev][Thread Next] - [Date Index][Thread Index][Author Index]

AO40: update



Dear Folks,

recovering of AO-40 continues and some housekeeping tasks were
performed by the command stations to improve and stabilize 
the situation. In addition, new software routines were successfully
loaded to restore the BCR-Service and other housekeeping functions.
The BCR's (Battery Charge Regulator) also work fine and we have
a positive power budget with happy batteries..
A quick look at the LIU instrumentation showed that the Helium
high pressure was essentially where we left it after the first
400N motor burn (173 bar).
It was found that a few temperature sensors have failed and 
also some of the current sensors are showing wrong values.
The sun sensors seem to work fine and showed a solar angle of 31 degs,
close to predicted. The spin rate needs some recalibrating. The
real spin rate appears to be in the range of 16 RPM (as measured
by the doppler wobbling on the S-Band beacon). This is a substantial
increase compared to the 9 RPM before the incident. This could be
due to a leak, although the difference in mass (m1/m2 = w2/w1) is
more than the available liquid fuel, thus the way how it left 
the spacecraft increased the spin too. Clearly, we need more time 
to analyze and understand what has happened here.
Basically it seems that the spacecraft is now fully under control
again. In the next days some more software will be loaded and
the various uplinks will be verified before we will try to turn
the 2m TX on again. However, there are no indications that the 
TX has failed, but we will make sure not to loose communication
again..  So for the next days the spacecraft will continue to
be transmitting on S-Band only..

73s Peter DB2OS for the whole command team

----
Via the amsat-bb mailing list at AMSAT.ORG courtesy of AMSAT-NA.
To unsubscribe, send "unsubscribe amsat-bb" to Majordomo@amsat.org



AMSAT Top AMSAT Home